# Flow matching for reaction pathway generation

### v3

@@ 5.2.4 Object-aware Message passing
Message passing is implemented with object-aware equivariance, where the model prediction is equivariant w.r.t. the relative rotation and translation of various molecules in a system [17].
@@ 5.2.5 Readout by an equivariant transformer
To efficiently capture the intricate interactions between atoms both within and across molecules, we employ the equivariant transformer (ET) [40] to update the scalar and vector features obtained in the previous section. The ET architecture preserves object-aware SE(3) symmetry, ensuring consistent treatment of rotations and translations, while effectively modeling complex geometric relationships. It exhibits strong scalability across large and diverse datasets, showing substantial performance gains when trained on extended molecular databases, as demonstrated by our results in the main text. This advantage is further reinforced by its computational efficiency, where the ET trains approximately 3–4× faster per epoch than existing architectures such as React-OT.
@@ 5.3 Model training
We train the model using the AdamW optimizer [41]. The learning rate is initialized at $1\times 10^{-4}$ and linearly warmed up to $5\times 10^{-4}$ over the first 10 epochs according to
@@ 5.4 Evaluation
IRC performed to verify the optimized TS is carried out using the Psi4 package [43]. The resulting endpoints of the IRC integration are further optimized to energy minimums. The comparison of IRC endpoints and input reactants and products is based on the adjacency matrix computed by the TAFFI package [29].
@@
For each reaction entry in the Transition1x test set, we extract the reactant and generate 240 corresponding product candidates. All generated samples are then combined, and unique reactants are identified. For each unique reactant, we compile the list of associated products and rank them based on their occurrence frequency. Molecular comparisons are performed using adjacency matrices computed with the TAFFI package [29].

### v4

@@ 5.2.4 Object-aware Message passing
Message passing is implemented with object-aware equivariance, where the model prediction is equivariant w.r.t. the relative rotation and translation of various molecules in a system [duan2023accurate].
@@ 5.2.5 Readout by an equivariant transformer
To efficiently capture the intricate interactions between atoms both within and across molecules, we employ the equivariant transformer (ET) [jiao2024equivariant] to update the scalar and vector features obtained in the previous section. The ET architecture preserves object-aware SE(3) symmetry, ensuring consistent treatment of rotations and translations, while effectively modeling complex geometric relationships. It exhibits strong scalability across large and diverse datasets, showing substantial performance gains when trained on extended molecular databases, as demonstrated by our results in the main text. This advantage is further reinforced by its computational efficiency, where the ET trains approximately 3–4× faster per epoch than existing architectures such as React-OT.
@@ 5.3 Model training
We train the model using the AdamW optimizer [loshchilov2017decoupled]. The learning rate is initialized at $1\times 10^{-4}$ and linearly warmed up to $5\times 10^{-4}$ over the first 10 epochs according to
@@ 5.4 Evaluation
IRC performed to verify the optimized TS is carried out using the Psi4 package [smith2020psi4]. The resulting endpoints of the IRC integration are further optimized to energy minimums. The comparison of IRC endpoints and input reactants and products is based on the adjacency matrix computed by the TAFFI package [seo2021topology].
@@
For each reaction entry in the Transition1x test set, we extract the reactant and generate 240 corresponding product candidates. All generated samples are then combined, and unique reactants are identified. For each unique reactant, we compile the list of associated products and rank them based on their occurrence frequency. Molecular comparisons are performed using adjacency matrices computed with the TAFFI package [seo2021topology].